Timetable Tuesday, September 10

Time  Room LG 1, HS 1  Room LG 1, HS 2  Room LG 1, HS 304

09.00 – 09.45 Plenary Talk (Chair: Andrzej Cegielski) Zdzisław Denkowski: On sensitivity of optimal solutions to control problems for hemivariational inequalities

09.50 – 10.20 (Chair: Vladimir Veliov) Danuta Przeworska-Rolewicz: A generalization of Lagrange, Du Bois-Reymond and Legendre lemmas using the algebraic analysis theory (Chair: Jacek Gondzio) Mauricio Solar: A parallel genetic scheduling algorithm

10.20 – 10.50 Andrej V. Dmitruk: An approximation theorem for nonlinear control systems with sliding modes and terminal equality constraints Ivo Nowak: Lagrangian deformation heuristics for solving MINLP’s

10.50 – 11.15  Coffee break

11.15 – 12.00 Plenary talk (Chair: Frédéric Bonnans) Ekkehard W. Sachs: Applications and efficient implementations of Gauss-Newton methods

12.05 – 12.35 (Chair: Bernd Kummer) Peter Fusek: Isolated zeros of Lipschitzian metrically regular functions (Chair: Valentin Gorokhovik) Christian Nagel: A trust-region smoothing-type method for semidefinite programs (Chair: Martin Gugat) Anatolij S. Antipin: Extraproximal method for solving a non-zero-sum twoperson game

12.35 – 13.05 Diethard Klatte: Constrained minima and Lipschitzian penalties Johannes Jahn: Connections between semidefinite programming and vector optimization Ścibór Sobieski: A new approach to the minimax problem

13.05 – 14.30  Lunch

14.30 - 15.15 Plenary talk (Chair: Diethard Klatte) Lionel Thibault: Regularity and proxregularity in variational analysis

Timetable Tuesday, September 10 (continued)

Time  Room LG 1, HS 1  Room LG 1, HS 2  Room LG 1, HS 304
15.15 – 15.45  Coffee break

15.45 – 16.15 Minisymposium (Chair: Ursula Felgenhauer) Kazimierz Malanowski: Stability and sensitivity analysis for state constrained optimal control. Some unsolved problems (Chair: Jiri Outrata) Xuan Duc Ha Truong: Ekeland’s variational principle for set-valued maps (Chair: Rainer Tichatschke) Marco A. López: Extended active constraints and some generic results in linear semi-infinite programming

16.15 – 16.45 Bernard Rousselet: Design sensitivity of controlled structures Darinka Dentcheva: Characterization of multifunctions by selections Armin Hoffmann: Coarse approach to generalized semi-infinite programming problems via global optimization

16.50 – 17.20 Minisymposium (Chair: Ursula Felgenhauer) Helmut Maurer: Sensitivity analysis of bang-bang control problems with applications to nonlinear laser optics (Chair: Christian Grossmann) Radoslaw Pytlak: The method of shortest residuals for large scale problems (Chair: Florian Jarre) Alexander Kruger: Some extended optimality notions are equivalent

17.20 – 17.50 Olga Kostyukova: Parametric optimal control problems with varying order of state constraints Andreas Fischer: The use of error bounds for analyzing and improving local behavior of algorithms Michael L. Flegel: First order optimality conditions for mathematical programs with equilibrium constraints

17.50 – 18.20 René Henrion: Point conditions for local error bounds, calmness and linear conditioning Andreas Grothey: A second order trust region bundle method for nonconvex nonsmooth optimization Andrzej Chmura: Factor-analysis of extremal problems

18.20 – 18.50 Alexander Ioffe: Stability estimates for regularity properties of nonlinear operators and set-valued maps François Courty: Trust region and one-shot approach to optimal control problems